I can't seem to get this one:
Which of the ordered pairs
(3,1), (0,-4), (-4,0), (-3,-7)
are solutions for the eqation x-y=4?
A) (-4,0) and (-3,-7)
B) (3,1) and (-4,0)
C) (0,-4),(-4,0), and (-3,-7)
D) (0,-4) and (-3,-7)

Answer by checkley71(8403)

Answer is D)(0,-4) & (-3,-7)
proof.
x-y=4 or x-(-4)=4 or 4=4
-3-(-7)=4 or -3+7=4 or 4=4